A freight train left a station and traveled at 30 mph. Two hours later, an express train left the same station and traveled in the same direction at 50 mph (on a different track). In how many hours did the express overtake the freight train?
Answer by ankor@dixie-net.com(22740)

A freight train left a station and traveled at 30 mph. Two hours later, an express train left the same station and traveled in the same direction at 50 mph (on a different track). In how many hours did the express overtake the freight train?
:
Let t = time it took the express to overtake the freight
:
Then (t+2) time the freight was traveling when it was overtaken
:
We know that when the express overtakes the freight, they will have traveled the same distance. Another words; Express dist = Freight dist
:
distance = speed * time:
:
A simple distance equation
Express dist = freight dist; (as we said before).
50t = 30(t+2)
50t = 30t + 60
50t - 30t = 60
20t = 60
t = 60/20
t = 3 hrs for the express to overtake the freight
:
:
Check our solution using distance:
50 * 3 = 30 * (3 + 2)
150 = 150
:
Make sense to you?